# Sina Baharlouei

### 06 Dec 2023 v1

@@ Appendix F Formal Statement of Theorem 2 and Proof
$G_{\mathcal{B}}=\frac{1}{|\mathcal{B}|}\sum_{(x_{i},y_{i})\in\mathcal{B}}\left[\nabla_{{\bm{\theta}}}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}{\mathbf{A}}_{jk}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)-f^{*}({\mathbf{A}}_{jk})\pi_{k}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\right]$
@@
$\left[\nabla_{{\bm{\theta}}}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}{\mathbf{A}}_{jk}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)-f^{*}({\mathbf{A}}_{jk})\pi_{k}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\right]$ (28)
@@
$\|{\mathbf{A}}_{jk}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)\|_{2}\leq L{\mathbf{A}}_{jk}$ (30)
$\|\pi_{k}f^{*}({\mathbf{A}}_{jk})\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\|_{2}\leq\pi_{k}Lf^{*}({\mathbf{A}}_{jk})$ (31)
These terms are bounded if ${\mathbf{A}}_{jk}$ is bounded and $f^{*}({\mathbf{A}}_{jk})$ is bounded for any ${\mathbf{A}}_{jk}$. This holds true for all $f$-divergences given assumptions 3 and 4. To see why, it suffices to find the optimal solution of each $f$-divergence by setting the gradient zero with respect to ${\mathbf{A}}_{jk}$. In all cases, the solution is a combination of $\mathbb{P}_{s_{k}}$ and $\mathbb{P}_{\hat{y}_{j}}$ terms that are non-zero and bounded (by assumptions 3 and 4). Since each term is bounded in (28), the expectation of the squared norm is also bounded, according to Lemma 23.
Finally, given that the estimator is unbiased, and the variance is bounded (Assumption 4.1 holds in Lin et al. [2020]), the two-time-scale stochastic gradient descent-ascent algorithm (which is Algorithm 1) finds an $\epsilon$-stationary solution of the Problem in $\mathcal{O}(\frac{1}{\epsilon^{8}})$ according to Theorem 4.9 in Lin et al. [2020].
Remark 24
For the case of strongly convex $f$-divergence (e.g. $\chi^{2}$ divergence), the convergence rate of $\mathcal{O}(\kappa^{3}\epsilon^{-4})$ can be obtained (Theorem 4.5 in Lin et al. [2020]). Such an iteration complexity holds for the batch size of $\mathcal{O}(\frac{\kappa\sigma^{2}}{\epsilon^{2}})$. If the batch size is as small as one, the rate will be $\mathcal{O}(\kappa^{3}\epsilon^{-5})$.
Remark 25
If the batch size is $n$ (deterministic), a rate of $\mathcal{O}(\epsilon^{-6})$ can be obtained according to Theorem 4.8 in Lin et al. [2020]. Note that it does not necessarily translate to a better runtime compared to the stochastic case. Because the per iteration cost of evaluating the gradient for $n$ data points can be much higher than evaluating on just $1$ (or a small number of) data points.
@@ Appendix G A Faster (But Double-Loop) First-order Optimization Algorithm for Optimizing ($f$-FERM)
We apply SREDA [Luo et al., 2020] to find an $\epsilon$ stationary solution of Problem ($f$-FERM). Note that SREDA works for non-convex strongly-concave min-max problems. We can directly apply the algorithm when $f$ is the $\chi^{2}$-divergence. In the cases that the function is concave but not strongly concave (e.g., KL divergence and Reverse KL), we first consider the following approximation:
@@
The maximization problem is an $\epsilon$-strongly concave problem. If we apply SREDA (see Algorithm 3 in Luo et al. [2020]), we find an $\epsilon$ stationary solution of Problem (32) in $\mathcal{O}(\kappa^{3}\epsilon^{-3})$ where $\kappa=\frac{L}{\mu}$ is the condition number. In our case, $\mu$, the strong concavity modulus can be set to the desired accuracy $\epsilon$ so that solving the approximate problem (32) leads to an approximate stationary point of the original problem. Therefore, the rate of convergence will be $\mathcal{O}(\epsilon^{-6})$. Note that applying SREDA finds an $\epsilon$ stationary solution of Problem (32). Similar to Theorem 3, since the added regularization term is small enough, the obtained solution is a $\mathcal{O}(\epsilon)$-stationary solution of the original problem ($f$-FERM). An important note is that the SREDA algorithm (line 13 in Algorithm 3 [Luo et al., 2020]) has a nested loop compared to the SGDA algorithm proposed in Algorithm 1. Therefore, the $\mathcal{O}(\epsilon^{-6})$ iteration complexity bound does not necessarily translate to an improved runtime in practice. Algorithm 2 describes SREDA applied to Problem (32). For the simplicity of the presentation, define the summation argument over $n$ data points as $\phi({\mathbf{x}}_{i},y_{i},s_{i},{\bm{\theta}},{\mathbf{A}})$.
@@
1:Input: periods $q$, m ¿ 0, Number of iterations T, step-size $\eta_{\bm{\theta}}$, fairness parameter $\lambda\geq 0,$ iteration number $T$, Batchsizes $S$ and $R$.
@@
The ConcaveMaximizer module is described in Algorithm 4 in Luo et al. [2020].
@@ Appendix H A First-order Optimization Algorithm for Optimizing (10)
The proof of the theorem is similar to Theorem 2 as the objective function is non-convex concave.
One can obtain faster algorithms under additional assumptions. For example, if the set for $\theta$ is assumed to be compact (e.g., we restrict the norm of the weight of the gradient), then we can accelerate the algorithm to $O(\epsilon^{-6})$, see Rafique et al. [2022]. Moreover, if we consider full batch sizes, we can utilize Algorithm 2 in Ostrovskii et al. [2021a]. This will give you the rate of convergence of $O(\epsilon^{-2})$ (Theorem 5.2).
@@ Appendix I Proof of Equation (12)
In this section, we perform ($f$-FERM), [Hardt et al., 2016][Mary et al., 2019], and [Baharlouei et al., 2020] to COMPAS  [^2] https://www.kaggle.com/datasets/danofer/compass and German Credit datasets  [^3] https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data. In the experiment on COMPAS, we use equality of opportunity as the measure of fairness violation, while in the German Credit dataset experiment, we use equalized odds. The results show that ($f$-FERM) is significantly better than other approaches regarding the accuracy-fairness tradeoff. The batch size is equal to $64$ for all methods.

### 07 Apr 2024 v2

@@ Appendix E Derivation of Closed-Form Expressions for Unbiased Gradient Estimators of $f$-Divergences
$G_{\mathcal{B}}=\frac{1}{|\mathcal{B}|}\sum_{(x_{i},y_{i})\in\mathcal{B}}\left[\nabla_{{\bm{\theta}}}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}A_{jk}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)-f^{*}(A_{jk})\pi_{k}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\right]$
@@
$\left[\nabla_{{\bm{\theta}}}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}A_{jk}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)-f^{*}(A_{jk})\pi_{k}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\right]$ (28)
@@
$\|A_{jk}\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)\|_{2}\leq LA_{jk}$ (30)
$\|\pi_{k}f^{*}(A_{jk})\nabla_{{\bm{\theta}}}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\|_{2}\leq\pi_{k}Lf^{*}(A_{jk})$ (31)
These terms are bounded if $A_{jk}$ is bounded and $f^{*}(A_{jk})$ is bounded for any $A_{jk}$. This holds true for all $f$-divergences given assumptions 3 and 4. To see why, it suffices to find the optimal solution of each $f$-divergence by setting the gradient zero with respect to $A_{jk}$. In all cases, the solution is a combination of $\mathbb{P}_{s_{k}}$ and $\mathbb{P}_{\hat{y}_{j}}$ terms that are non-zero and bounded (by assumptions 3 and 4). Since each term is bounded in (28), the expectation of the squared norm is also bounded, according to Lemma F.6.
Finally, given that the estimator is unbiased, and the variance is bounded (Assumption 4.1 holds in Lin et al. (2020)), the two-time-scale stochastic gradient descent-ascent algorithm (which is Algorithm 1) finds an $\epsilon$-stationary solution of the Problem in $\mathcal{O}(\frac{1}{\epsilon^{8}})$ according to Theorem 4.9 in Lin et al. (2020). ∎
Remark F.7.
For the case of strongly convex $f$-divergence (e.g. $\chi^{2}$ divergence), the convergence rate of $\mathcal{O}(\kappa^{3}\epsilon^{-4})$ can be obtained (Theorem 4.5 in Lin et al. (2020)). Such an iteration complexity holds for the batch size of $\mathcal{O}(\frac{\kappa\sigma^{2}}{\epsilon^{2}})$. If the batch size is as small as one, the rate will be $\mathcal{O}(\kappa^{3}\epsilon^{-5})$.
Remark F.8.
If the batch size is $n$ (deterministic), a rate of $\mathcal{O}(\epsilon^{-6})$ can be obtained according to Theorem 4.8 in Lin et al. (2020). Note that it does not necessarily translate to a better runtime than the stochastic case. Because the per iteration cost of evaluating the gradient for $n$ data points can be much higher than evaluating on just $1$ (or a small number of) data points.
@@ Appendix G A Faster (But Double-Loop) First-order Optimization Algorithm for Optimizing ($f$-FERM)
We apply SREDA (Luo et al., 2020) to find an $\epsilon$ stationary solution of Problem ($f$-FERM). Note that SREDA works for non-convex strongly-concave min-max problems. We can directly apply the algorithm when $f$ is the $\chi^{2}$-divergence. In the cases that the function is concave but not strongly concave (e.g., KL divergence and Reverse KL), we first consider the following approximation:
@@
The maximization problem is an $\epsilon$-strongly concave problem. If we apply SREDA (see Algorithm 3 in Luo et al. (2020)), we find an $\epsilon$ stationary solution of Problem (32) in $\mathcal{O}(\kappa^{3}\epsilon^{-3})$ where $\kappa=\frac{L}{\mu}$ is the condition number. In our case, $\mu$, the strong concavity modulus can be set to the desired accuracy $\epsilon$ so that solving the approximate problem (32) leads to an approximate stationary point of the original problem. Therefore, the rate of convergence will be $\mathcal{O}(\epsilon^{-6})$. Note that applying SREDA finds an $\epsilon$ stationary solution of Problem (32). Similar to Theorem 3.1, since the added regularization term is small enough, the obtained solution is a $\mathcal{O}(\epsilon)$-stationary solution of the original problem ($f$-FERM). An important note is that the SREDA algorithm (line 13 in Algorithm 3 (Luo et al., 2020)) has a nested loop compared to the SGDA algorithm proposed in Algorithm 1. Therefore, the $\mathcal{O}(\epsilon^{-6})$ iteration complexity bound does not necessarily translate to an improved runtime in practice. Algorithm 2 describes SREDA applied to Problem (32). For the simplicity of the presentation, define the summation argument over $n$ data points as $\phi({\mathbf{x}}_{i},y_{i},s_{i},{\bm{\theta}},{\mathbf{A}})$.
@@
1:Input: periods $q$, m > 0, Number of iterations T, step-size $\eta_{\bm{\theta}}$, fairness parameter $\lambda\geq 0,$ iteration number $T$, Batchsizes $S$ and $R$.
@@
The ConcaveMaximizer module is described in Algorithm 4 in Luo et al. (2020).
@@ Appendix H A First-order Optimization Algorithm for Optimizing (10)
The proof of the theorem is similar to Theorem 2.2 as the objective function is non-convex concave.
One can obtain faster algorithms under additional assumptions. For example, if the set for $\theta$ is assumed to be compact (e.g., we restrict the norm of the weight of the gradient), then we can accelerate the algorithm to $O(\epsilon^{-6})$, see Rafique et al. (2022). Moreover, if we consider full batch sizes, we can utilize Algorithm 2 in Ostrovskii et al. (2021a). This will give you the rate of convergence of $O(\epsilon^{-2})$ (Theorem 5.2).
@@ Appendix I Proof of Equation (12)
In this section, we perform ($f$-FERM), (Hardt et al., 2016)(Mary et al., 2019), and (Baharlouei et al., 2020) to COMPAS  [^2] https://www.kaggle.com/datasets/danofer/compass and German Credit datasets  [^3] https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data. In the experiment on COMPAS, we use equality of opportunity as the measure of fairness violation, while in the German Credit dataset experiment, we use equalized odds. The results show that ($f$-FERM) is significantly better than other approaches regarding the accuracy-fairness tradeoff. The batch size is equal to $64$ for all methods.